# Adaptive Self-improvement LLM Agentic System for ML Library Development

### v1

@@ 4.4 Structual intermediate representation
Good interfaces can improve the performance of agentic coding systems (Yang et al., 2024; Wei et al., 2024). We borrow the intermediate representation (IR) technique from compiler literature (Lattner et al., 2021) and use a structural IR to unify the interfaces of our agentic system. Specifically, the structural IR is used as the interface between users and the agentic system and as the interface between LLM agents and the code generator within the system.
@@ 5.1 Metric
[…]
5.2 Construct the benchmark suite
@@ 6 Experiments
Detailed experimental settings are in Section A.3. We also benchmark the tasks with OpenAI-o1 in Table 4 but do not include it in the following experiments to control for test-time compute. All prompts are formatted in YAML because structural prompts generally benefit (He et al., 2024).
@@ 6.1 Analysis of adaptive self-improvement learning
This section uses the agentic system organization described in Section 4.1 for the best possible base learning capability.
The hard-only examples improve the performance more effectively than examples mixed with easier ones. As shown in Figure 7, the Pareto optimal is composed of hard examples (denoted by “H”) for all three models. Moreover, hard examples bring the most significant improvement along the learning curve. In some cases, fewer hard examples may perform better than more examples mixed with easier examples. For example, Iter 3 has better performance than Iter 2 for gpt-4o.
@@
Adaptive granularity affects token efficiency and peak performance. If the adaptive granularity $m$ is too small, then easy examples might dilute the difficulty of training data. If $m$ is too large, then exploration steps might be too conservative and thus waste input tokens. Therefore, there is a sweet spot that balances the training data difficulty and cost of input tokens. As shown in Figure 8, $m=3$ is that spot. $m=3$ saves 1.07×1.07$\times$1.07 × tokens over $m=4$ while maintaining performance. Additionally, $m=3$ improves the performance by 1.5×1.5$\times$1.5 × at a similar input token cost when compared to $m=1$.
@@ 6.2 Ablation study of agentic system organization
Agentic system can discover non-trivial STeP programs. Surprisingly, the agentic system composes attention operators with over 50% Pass@1 as shown in Figure 13. That means the agentic system can discover online softmax (Milakov & Gimelshein, 2018) and memory-free streaming attention (Sohn et al., 2024b) with specialized functions provided in Figure 16, which is considered challenging for ordinary programmers.
@@
Table 1: Analysis of the correlation between semantic diversity of answers and the performance. “-WS” means “without structural IR”. Higher values indicate higher semantic diversity.
The guardian agent can correct affine type errors but might also corrupt correct answers output by the proposer agent.
@@ 7.1 Self-improvement learning for LLMs
Self-improvement learning for LLMs typically involves two stages: scoring generated samples (trajectories) and incorporating those samples to enhance the model. Scoring can be achieved through human labeling (Cobbe et al., 2021; Lightman et al., 2023) or through automated methods such as verifiers and heuristics (Wang et al., 2024; Singh et al., 2023). Our method stands out in this context by utilizing AST analysis, offering a more interpretable approach to scoring. When it comes to incorporating samples, models may rely on retrieval (Zhao et al., 2024; Park et al., 2023), reflection (Shinn et al., 2024; Liu et al., 2023), or reward feedback (Opsahl-Ong et al., 2024; Fernando et al., 2023). Our method introduces a new mechanism in this stage by adaptively extending and prioritizing high-scoring samples.
@@ 7.2 Agentic system organization for specialized tasks
Task-specific organization has proven effective in enhancing the performance of agentic systems across diverse coding tasks (Zhang et al., 2024b; Fang et al., 2024; Guan et al., 2024). We adopt an agentic system organization specifically for ML library development using an ASPL. Such domain knowledge can be further augmented with automatic agentic system design tools (Khattab et al., 2023; Hu et al., 2024). Furthermore, well-designed interfaces between agents, tools, and other agents have been shown to improve performance (Schick et al., 2023; Yang et al., 2024; Wu et al., 2023), and a structural IR enables these interfaces to be highly task-aligned in our system.
@@ 8 Conclusion
ML library development using ASPLs is a critical component of the ML ecosystem, but it remains poorly automated. To address this limitation, we co-design the learning process and agentic system around a central objective: enabling complex reasoning with limited data. Our methods simultaneously implement non-trivial ML operators and produce a self-improving agent. Consequently, our system provides not only support to experts in developing ML libraries but also valuable resources for other systems.
@@ A.1 STeP introduction
0:  $X$: [m,n] of Buffer(k), $N$: [m,n] of Multihot(e), $G$: [m,n] of Buffer(e)
@@
1:  $S_{0}$ = Zip($X$, $G$)
2:  $S_{1}^{0}$, $S_{1}^{1}$ = Copy($N$)
3:  $S_{2}$ = Partition(2, $S_{0}$, $S_{1}^{0}$)
4:  $S_{3}$ = [Map(expert0, $S_{2}$[0]), Map(expert1, $S_{2}$[1])]
5:  $S_{4}$ = Merge(weightedsum, $S_{3}$, $S_{1}^{1}$)
@@
[…]
@@ A.2 Benchmark details
As shown in Figure 3, Group 4 tasks have the same operator: $softmax(S)\cdot V$ where $S$ equals $QK^{T}$. They differ in external functions as shown in Figure 16. Group 4 can use RDA’s on-chip fusion to compose scale-dot-product attention with Group 5. Group 7 also differs in external functions. Group 5 contains three dataflow orders: inner-product(”$mnk,mdk\rightarrow mnd$”), row-wise(”$mnk,mkd\rightarrow mnd$”), and outer-product(”$mkn,mkd\rightarrow mnd$”). Group 6 contains GptJ and NeoX styles which differ in pairing even and odd or the first and second half positions (vLLM, 2023). Group 8 contains LayerNorm and RMSNorm.
@@ A.3 Experiment settings
In Section 6.1 we sample 64 times for each temperature of 0.4, 0.7, and 1.0, recording the best result. In Section 6.2, we sample 64 times at temperature 0.7 on Claude 3.5 Sonnet to control variables. Four models are: claude-3-5-sonnet-20241022 of Anthropic API (Claude 3.5 Sonnet), gpt-4o-2024-11-20 of OpenAI API (GPT-4o), deepseek-chat of DeepSeek API (DeepSeek-V3), and Meta-Llama-3-1-405B-Instruct-Turbo of TogetherAI API (Llama 3.1-405B). Maximum output tokens are set as 1024 and the seed for GPT-4o is 42.
A.4 Extra results
@@
As shown in Table 4, OpenAI-o1 achieves similar performance at the cost of more tokens than a single Claude-3-5-Sonnet. This observation aligns with previous findings that scaling pretraining is preferable over inference for challenging tasks (Snell et al., 2024). Meanwhile, a single Claude-3-5-Sonnet proposer can finish more tasks than the OpenAI-o1 using fewer tokens. Table 3 contains all abbreviations.
@@
A.5 Additional explanation
@@
A.6 Code and prompt examples
@@
# Task 0 # MaxSum def apply(state, input): $m_{t},l_{t},\boldsymbol{o_{t}}$ = state $s_{t},\boldsymbol{v_{t}}$ = input $m_{t+1}=max(m_{t},s_{t})$ $l^{\prime}_{t}=l_{t}*e^{(m_{t}-m_{t+1})}$ $p_{t}=e^{(s_{t}-m_{t+1})}$ $l_{t+1}=p_{t}+l^{\prime}_{t}$ $\boldsymbol{o_{t+1}}=\frac{1}{l_{t+1}}(l^{\prime}_{t}*\boldsymbol{o_{t}}+p_{t}*\boldsymbol{v_{t}})$ return $(m_{t+1},l_{t+1},\boldsymbol{o_{t+1}})$ def init(): return $(-\infty,0,\boldsymbol{0})$ GetThrid def apply(input): return input[2] # Task 1 # ExpMaxDiff def apply(state, input): $m_{t},e_{t},d_{t}$ = state $s_{t},$ = input $m_{t+1}=max(m_{t},s_{t})$ $\Delta m=m_{t}-m_{t+1}$ $e_{t+1}=e^{(s_{t}-m_{t+1})}$ $d_{t+1}=e^{\Delta m}$ return $(m_{t+1},e_{t+1},d_{t+1})$ def init(): return $(-\infty,0,0)$ DivSum def apply(state, input): $\boldsymbol{v_{t}},e_{t},d_{t}$ = input $l_{t},\boldsymbol{o_{t}}$ = state $l^{\prime}_{t}=d_{t}*l_{t}$ $l_{t+1}=e_{t}+l^{\prime}_{t}$ $o_{t+1}=\frac{1}{l_{t+1}}(l^{\prime}_{t}*\boldsymbol{o_{t}}+e_{t}*\boldsymbol{v_{t}})$ return $(l_{t+1},o_{t+1})$ def init(): return $(0,\boldsymbol{0})$ GetSecondThrid def apply(input): return input[1],input[2] # GetSecond def apply(input): return input[1] # Task 2 # ExpMaxDiff def apply(state, input): $m_{t},e_{t},d_{t}$ = state $s_{t},$ = input $m_{t+1}=max(m_{t},s_{t})$ $\Delta m=m_{t}-m_{t+1}$ $e_{t+1}=e^{(s_{t}-m_{t+1})}$ $d_{t+1}=e^{\Delta m}$ return $(m_{t+1},e_{t+1},d_{t+1})$ def init(): return $(-\infty,0,0)$ GetSecondThrid def apply(input): return input[1],input[2] # WeightedSumSingle def apply(state, input): $e_{t},d_{t}$ = input $r_{t}$ = state return $(r_{t}*d_{t}+e_{t})$ def init(): return 00 # WeightedSumDouble def apply(state, input): $\boldsymbol{v_{t}},e_{t},d_{t}$ = input return $(state*d_{t}+e_{t}*\boldsymbol{v_{t}})$ def init(): return $\boldsymbol{0}$ Div def apply(input): $r_{t},\boldsymbol{l_{t}}=input$ return $\frac{\boldsymbol{l_{t}}}{r_{t}}$

### v2

@@ 4.4 Structual intermediate representation
Good interfaces can improve the performance of agentic coding systems (Yang et al., 2024; Wei et al., 2024). Therefore, we are inspired by the intermediate representation (IR) technique from compiler literature (Lattner et al., 2021; Vasilache et al., 2022) and use a structural IR to unify the interfaces of our agentic system. Specifically, this structural IR is the interface between users and the agentic system and between LLM agents and the code generator within the system.
@@ 5.1 Metric
$\text{pass@k}\coloneqq\frac{1}{T}\sum_{i=1}^{T}\left[1-\frac{\binom{n-c_{i}}{k}}{\binom{n}{k}}\right$] (1)
5.2 Benchmark Construction
@@ 6 Experiments
Detailed experimental settings are in Section A.3. We also benchmark the tasks with OpenAI-o1 in Table 4 but do not include it in the following experiments to control for test-time compute. As described in Section 4.4, all prompts are formatted in YAML because structural prompts generally benefit (He et al., 2024).
@@ 6.1 Analysis of adaptive self-improvement learning
This section uses the agentic system organization described in Section 4.1 for the best possible base learning capability. Our evaluation provides the following insights:
The hard-only examples improve performance more effectively than examples mixed with easier ones. As shown in Figure 7, the Pareto optimal is composed of hard examples (denoted by “H”) for all three models. Moreover, hard examples bring the most significant improvement along the learning curve. In some cases, fewer hard examples may perform better than more examples mixed with easier examples. For example, Iter 3 has better performance than Iter 2 for gpt-4o.
@@
Adaptive granularity affects token efficiency and peak performance. If the adaptive granularity $m$ is too small, then easy examples might dilute the difficulty of training data. If $m$ is too large, then exploration steps might be too conservative and thus waste input tokens. Therefore, there is a sweet spot that balances the training data difficulty and cost of input tokens. As shown in Figure 8, $m=3$ is that spot. $m=3$ saves $1.07\times$ tokens over $m=4$ while maintaining performance. Additionally, $m=3$ improves the performance by $1.5\times$ at a similar input token cost when compared to $m=1$.
@@ 6.2 Ablation study of agentic system organization
Table 1: Analysis of the correlation between semantic diversity of answers and the performance. “-w/o-IR” means “without structural IR”. Higher values indicate higher semantic diversity.
The agentic system can discover non-trivial STeP programs. Surprisingly, the agentic system composes attention operators with over 50% Pass@1 as shown in Figure 13. That means the agentic system can discover online softmax (Milakov & Gimelshein, 2018) and memory-free streaming attention (Sohn et al., 2024b) with specialized functions provided in Section A.7, which is considered challenging for ordinary programmers.
@@
The guardian agent can correct affine type errors but might corrupt the correct answers output by the proposer agent.
@@ 7 Related Work
This work introduces a self-improvement agentic system for ASPL code generation, enhancing LLM effectiveness through an adaptive learning algorithm and deliberate agentic system design. Accordingly, we focus the related work on approaches that leverage LLMs for self-improvement and agentic systems on specialized tasks.
@@ 7.1 Self-improvement learning for LLMs
Self-improvement learning for LLMs typically involves two stages: scoring generated samples (trajectories) and incorporating those samples to enhance the model. Scoring can be achieved through human labeling (Cobbe et al., 2021; Lightman et al., 2023) or through automated methods such as verifiers and heuristics (Wang et al., 2024; Singh et al., 2023). Our method stands out in this context by utilizing AST analysis, offering a more interpretable approach to scoring. When it comes to incorporating samples, models may rely on retrieval (Zhao et al., 2024; Park et al., 2023), reflection (Shinn et al., 2024; Liu et al., 2023; Yao et al., 2023), or reward feedback (Opsahl-Ong et al., 2024; Fernando et al., 2023). Our method introduces a new mechanism in this stage by adaptively extending and prioritizing high-scoring samples.
@@ 7.2 Agentic system organization for specialized tasks
Task-specific organization has proven effective in enhancing the performance of agentic systems across diverse coding tasks (Zhang et al., 2024c; Fang et al., 2024; Guan et al., 2024). We adopt an agentic system organization specifically for ML library development using an ASPL. Such domain knowledge can be further augmented with automatic agentic system design tools (Khattab et al., 2023; Hu et al., 2024; Zhang et al., 2024a). Furthermore, well-designed interfaces between agents, tools, and other agents have been shown to improve performance (Schick et al., 2023; Yang et al., 2024; Wu et al., 2023), and a structural IR enables these interfaces to be highly task-aligned in our system.
@@ 8 Conclusion
ML library development using ASPLs is a critical component of the ML ecosystem, but it remains poorly automated. To address this limitation, we co-design the learning process and agentic system around a central objective: enabling complex reasoning with limited data. Our methods simultaneously implement non-trivial ML operators and produce a self-improving agent. Consequently, our system not only supports experts in developing ML libraries but also offers valuable resources for other systems. We recognize that the challenges of complex reasoning under limited data extend beyond this domain and envision applying similar self-improvement techniques with LLM agents to a broad class of tasks.
@@
Acknowledgement
We thank Vishnu Sarukkai, Haochen Shi, Xinhao Li, Gina Sohn, Nathan Zhang, Tian Zhao, Yu Sun, Anjiang Wei, Chun Deng, Qizheng Zhang, Rubens Lacouture, Fredrik Kjolstad and the anonymous reviewers for their feedback on this paper. This research was supported in part by the Pervasive Parallelism Lab Affiliates Program Fund. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the aforementioned funding agencies.
@@ A.1 STeP introduction
0:$X$: [m,n] of Buffer(k), NN: [m,n] of Multihot(e), $G$: [m,n] of Buffer(e)
@@
[…]
@@
As shown in Figure 10, Copy duplicates a stream for the affine type constraint. Zip combines two streams of values into a stream of tuples. Map applies the function (fn) on each input value. Partition routes tokens of the data stream to experts assigned by the index stream ($S_{1}^{0}$). Merge accumulates tokens from experts. The accumulation of Merge is parameterized by fn and init where init initializes the state and fn updates the state with the input value. Partition and Merge are used in pairs, sharing the same index stream. $\langle\rangle$ represents the Tuple type of value tokens. Buffer, Multihot, and Scalar are also types of value tokens, parametrized by the generic data type like float and half, which is omitted for simplicity in the algorithm. Buffer and Multihot types are further parametrized by the shape in the paratheses. Shape manipulation primitives include Promote, Repeat and RepeatRefFigure 4(a) also shows two streams. S1 is a control token signaling the end of rank-1. 01 is a value token of multihot vector type served in index streams. (v0,v1) is a value token of type Tuple(Scalar, Reference) because the weight and input streams are composed of scalar and reference values, respectively.
@@ A.2 Benchmark details
As shown in Figure 3, Group 4 tasks have the same operator: $softmax(S)\cdot V$ where $S$ equals $QK^{T}$. They differ in external functions as shown in Section A.7. Group 4 can use RDA’s on-chip fusion to compose scale-dot-product attention with Group 5. Group 7 also differs in external functions. Group 5 contains three dataflow orders: inner-product($mnk,mdk\rightarrow mnd$”), row-wise($mnk,mkd\rightarrow mnd$”), and outer-product($mkn,mkd\rightarrow mnd$”). Group 6 contains GptJ and NeoX styles which differ in pairing even and odd or the first and second half positions (vLLM, 2023). Group 8 contains LayerNorm and RMSNorm.
@@ A.3 Experiment settings
In Section 6.1 we sample 64 times for each temperature of 0.4, 0.7, and 1.0, recording the best result. In Table 1, we sample 64 times at temperature 0.7 on Claude 3.5 Sonnet to control variables. Four models are: claude-3-5-sonnet-20241022 of Anthropic API (Claude 3.5 Sonnet), gpt-4o-2024-11-20 of OpenAI API (GPT-4o), deepseek-chat of DeepSeek API (DeepSeek-V3), and Meta-Llama-3-1-405B-Instruct-Turbo of TogetherAI API (Llama 3.1-405B). Maximum output tokens are set as 1024 and the seed for GPT-4o is 42.
A.4 Additional experiments
@@
As shown in Table 4, OpenAI-o1 achieves similar performance at the cost of more tokens than a single Claude-3-5-Sonnet. This observation aligns with previous findings that scaling pretraining is preferable over inference for challenging tasks (Snell et al., 2024). Meanwhile, a single Claude-3-5-Sonnet proposer can finish more tasks than the OpenAI-o1 using fewer tokens. Table 3 contains abbreviations for all tasks.
@@
Our system completed each task in under 10 minutes on average. In contrast, during our pilot study, a domain expert was unable to write a single program within 48 hours, as they had to do trial-and-error and accumulate experience sequentially. Our system, by comparison, can perform these explorations in parallel. In the future, collaboration with HCI researchers will enable more extensive experiments comparing time and effort required by human programmers versus the system, providing quantitative data on usability and cognitive load.
Differences in tool access and computational load might have influenced the outcomes. Therefore, we conducted an experiment that aligned both aspects.
For computation fairness, we matched the token count of the single model with the agent and self-improved models by resampling. All model variants (single, agent, self-improved) have access to the same verifier so it is fair; the difference lies in how the verifier is leveraged. Self-improved models incorporate it throughout the process, while others use it only at the end as a final judge. We chose Claude-3-5-Sonnet and GPT-4o as base models. As shown in Table 5, our agentic systems still perform better under this fair setting.
Pass@n Claude Sonnet Claude Sonnet GPT-4o GPT-4o
Single Agent & Self-improved Single Agent & Self-improved
From 0.73 0.73 0.23 0.23
To 0.77 0.96 0.38 0.81
Table 5: Performance comparison of single models vs agent & self-improved models with token count matching.
We conducted supervised finetuning (SFT) using GPT-4o. Table 6 shows that SFT can improve performance, but less than our self-improvement approach.
Since we do not know the exact SFT algorithm of OpenAI service for FLOPs matching, we tried our best to favor the SFT method. We began with the same 133 correct samples from all completed tasks used in the first iteration of self-improvement. Different from self-improvement which only picks 1 correct program per completed task, we picked all 133 programs to form the training dataset of SFT. We created three SFT datasets with varying prompt compositions:
  • 133 (base prompt+question+answer)
  • 133 (question+answer)
  • 17 (question+answers deduplicated via AST)
Each dataset was used to train a separate SFT model. After that, we sampled each model on all the uncompleted tasks.
Pass@n Finetuned Self-improved
From 0.35 0.35
To 0.62 0.81
Table 6: Performance comparison between finetuned and self-improved Models
We assessed code maintainability using two metrics: maintenance index without comments (MIwoc) and with comments (MI)  [^2] https://www.verifysoft.com/en_maintainability.html. The comment weight (MIcw) is defined as MI - MIwoc and falls in [0, 49); MI >> 85 indicates good maintainability. Using all correct programs from our best model (self-improved agentic Claude Sonnet), we recorded the top MIwoc and MI per task. The mean MIwoc is 102, MI is 149, and MIcw is 47—indicating well-commented, maintainable code.
As shown in Table 7, we used the same set of programs as the maintainability statistics to measure complexity. We measured these three metrics:
  • Lines of code: Counted via primitive calls (excluding comments/blank lines)
  • Shape transformations: Counted by use of Promote, Repeat, RepeatRef, and Flatten primitives
  • Specialized instructions: Counted as the number of specialized functions in task descriptions
Metric Min Max Mean
Lines of code 4 17 8.67
Shape transformations 0 6 1.13
Specialized instructions 2 7 3.68
Table 7: The assessment of complexity across all completed tasks.
A.5 Generalization
The proposed system can be extended to other scenarios that require complex reasoning with limited example data and well-defined evaluation metrics. We outline the general recipe below.
As shown in Figure 1, the agentic system organization is constructed in three main steps. First, system designers define the format of both the task and its expected output. Once the format is specified, the next step is to build a verifier for the task. With the format and verifier in place, designers can either use a single LLM or design LLM agents tailored to the domain—similar to how we handle the type constraints of the STeP language. After completing these three steps, the task can be handed over to our system, which will automatically carry out adaptive self-improvement learning.
The adaptive self-improvement learning system also exposes several tunable hyperparameters which will be helpful when the results are not satisfactory. The most direct control is the number of parallel sampling. Users can also adjust the adaptive granularity parameter m for experience stratification. Additionally, domain-specific filtering heuristics—such as the minimal code length heuristic used by us—can be incorporated to further guide the learning process.
Our approach has two parts: adaptive self-improvement learning and agentic system organization. The learning process is broadly applicable to other programming languages; the challenges lie in tailoring agentic systems to other languages.
Mainstream languages like CUDA, HIP, and CPU vector intrinsics exhibit global properties such as arbitrary memory access, data layout sensitivity, and side effects. Similarly, STeP enforces a global affine type constraint. Our framework addressed this using a guardian agent that detects and corrects affine type violations. This concept generalizes: domain-specific guardian agents can monitor and enforce global properties of various languages, adapting the STeP solution more broadly.
A second challenge is that LLMs may lean toward surface-level patterns in mainstream languages due to their existence in training data, potentially missing more optimal or novel transformations. As shown in Section 6.2, our structural IR can increase sample diversity and thus boost the LLM agentic system performance. Extending this, structural IRs and tailored code generators can guide LLMs toward more creative solutions beyond conventional patterns.
We also conducted an experiment on the AIME-2024 dataset  [^3] https://huggingface.co/datasets/Maxwell-Jia/AIME_2024 which contains 30 challenging problems from the American Invitational Mathematics Examination (AIME) 2024. We applied our adaptive self-improvement learning to the Claude-3-5-Sonnet base model and increased Pass@n from 0.50 to 0.67. This demonstrates the potential capabilities of our system on other tasks.
A.6 Additional explanation
@@
Existing auto-tuning code generation systems like Spiral (Puschel et al., 2005), Ansor (Zheng et al., 2020), and AKG (Zhao et al., 2021) rely on predefined optimization strategies and cost models. On the contrary, we employ LLMs to generate and refine code for machine learning libraries, particularly targeting ASPLs. We explore the shift towards utilizing the reasoning capabilities of LLMs for code generation and optimization, enabling the system to adapt and improve over time without explicit human intervention.
A.7 Code and prompt examples
@@
# Task 0 # MaxSum def apply(state, input): $m_{t},l_{t},\boldsymbol{o_{t}}$ = state $s_{t},\boldsymbol{v_{t}}$ = input $m_{t+1}=max(m_{t},s_{t})$ $l^{\prime}_{t}=l_{t}*e^{(m_{t}-m_{t+1})}$ $p_{t}=e^{(s_{t}-m_{t+1})}$ $l_{t+1}=p_{t}+l^{\prime}_{t}$ $\boldsymbol{o_{t+1}}=\frac{1}{l_{t+1}}(l^{\prime}_{t}*\boldsymbol{o_{t}}+p_{t}*\boldsymbol{v_{t}})$ return $(m_{t+1},l_{t+1},\boldsymbol{o_{t+1}})$ def init(): return $(-\infty,0,\boldsymbol{0})$ # GetThrid def apply(input): return input[2] # Task 1 # ExpMaxDiff def apply(state, input): $m_{t},e_{t},d_{t}$ = state $s_{t},$ = input $m_{t+1}=max(m_{t},s_{t})$ $\Delta m=m_{t}-m_{t+1}$ $e_{t+1}=e^{(s_{t}-m_{t+1})}$ $d_{t+1}=e^{\Delta m}$ return $(m_{t+1},e_{t+1},d_{t+1})$ def init(): return $(-\infty,0,0)$ # DivSum def apply(state, input): $\boldsymbol{v_{t}},e_{t},d_{t}$ = input $l_{t},\boldsymbol{o_{t}}$ = state $l^{\prime}_{t}=d_{t}*l_{t}$ $l_{t+1}=e_{t}+l^{\prime}_{t}$ $o_{t+1}=\frac{1}{l_{t+1}}(l^{\prime}_{t}*\boldsymbol{o_{t}}+e_{t}*\boldsymbol{v_{t}})$ return $(l_{t+1},o_{t+1})$ def init(): return $(0,\boldsymbol{0})$ # GetSecondThrid def apply(input): return input[1],input[2] # GetSecond def apply(input): return input[1] # Task 2 # ExpMaxDiff def apply(state, input): $m_{t},e_{t},d_{t}$ = state $s_{t},$ = input $m_{t+1}=max(m_{t},s_{t})$ $\Delta m=m_{t}-m_{t+1}$ $e_{t+1}=e^{(s_{t}-m_{t+1})}$ $d_{t+1}=e^{\Delta m}$ return $(m_{t+1},e_{t+1},d_{t+1})$ def init(): return $(-\infty,0,0)$ # GetSecondThrid def apply(input): return input[1],input[2] # WeightedSumSingle def apply(state, input): $e_{t},d_{t}$ = input $r_{t}$ = state return $(r_{t}*d_{t}+e_{t})$ def init(): return 0 # WeightedSumDouble def apply(state, input): $\boldsymbol{v_{t}},e_{t},d_{t}$ = input return $(state*d_{t}+e_{t}*\boldsymbol{v_{t}})$ def init(): return $\boldsymbol{0}$ # Div def apply(input): $r_{t},\boldsymbol{l_{t}}=input$ return $\frac{\boldsymbol{l_{t}}}{r_{t}}$